OpenStudy (anonymous):

Solve 3x2 + 18x + 15 = 0 by completing the square. 3(x + 3)2 = 6; x = −3 + square root 2, x = −3 − square root 2 3(x + 9)2 = 12; x = −11, x = −7 3(x + 3)2 = 12; x = −5, x = −1 3(x + 9)2 = 228; x = −9 + square root 76, x = −9 − square root 76

OpenStudy (anonymous):

@Hitaro9

OpenStudy (anonymous):

@flowmaster1776 @W.W.G.D @quickstudent @texaschic101 @TheSmartOne @Preetha @Pi123 @darkprince14 @UsukiDoll @vera_ewing @naledimunetsi @Samus_Aran @Madgirlwithabluebox @Mrs.Irwin

OpenStudy (anonymous):

YES PERSON

OpenStudy (usukidoll):

no mass tagging!

OpenStudy (anonymous):

Whaaaaaat? I really need help and I didn't know it wasn't allowed... Sorry!

OpenStudy (anonymous):

can you help though?

OpenStudy (usukidoll):

I got my own problems... if you know matlab that would be nice.

OpenStudy (anonymous):

okay thanks anyway

OpenStudy (madgirlwithabluebox):

Im not any good at math sorry :P

OpenStudy (anonymous):

okay :(

OpenStudy (anonymous):

darkprince?

OpenStudy (anonymous):

I'm thinking D but I checked other places and That's not what anyone else got...

OpenStudy (anonymous):

I just don't understand why

OpenStudy (anonymous):

guys...?

OpenStudy (darkprince14):

Completing the square means we want to transform ax2 + bx + c into the form a(x+d)2 + e=0 Where d =b/2a And e = c - (b^2 /4a)

OpenStudy (anonymous):

so 3(x^2 + 6x + 8) + 228

OpenStudy (anonymous):

that's what I got

OpenStudy (anonymous):

UsukiDoll you're still here?

OpenStudy (darkprince14):

Note that 3x^2 + 18x + 15=0 is of the form ax^2 + bx + c=0 From there, find a, b and c and then apply them to the equation i have provided earlier